# VQ-CTAP: Cross-Modal Fine-Grained Sequence Representation Learning for Speech Processing

### v1

@@ Abstract
Deep learning has brought significant improvements to the field of cross-modal representation learning. For tasks such as text-to-speech (TTS), voice conversion (VC), and automatic speech recognition (ASR), a cross-modal fine-grained (frame-level) sequence representation is desired, emphasizing the semantic content of the text modality while de-emphasizing the paralinguistic information of the speech modality. We propose a method called ”Vector Quantized Contrastive Token-Acoustic Pre-training (VQ-CTAP)”, which uses the cross-modal aligned sequence transcoder to bring text and speech into a joint multimodal space, learning how to connect text and speech at the frame level. The proposed VQ-CTAP is a paradigm for cross-modal sequence representation learning, offering a promising solution for fine-grained generation and recognition tasks in speech processing. The VQ-CTAP can be directly applied to VC and ASR tasks without fine-tuning or additional structures. We propose a sequence-aware semantic connector, which connects multiple frozen pre-trained modules for the TTS task, exhibiting a plug-and-play capability. We design a stepping optimization strategy to ensure effective model convergence by gradually injecting and adjusting the influence of various loss components. Furthermore, we propose a semantic-transfer-wise paralinguistic consistency loss to enhance representational capabilities, allowing the model to better generalize to unseen data and capture the nuances of paralinguistic information. In addition, VQ-CTAP achieves high-compression speech coding at a rate of 25Hz from 24kHz input waveforms, which is a 960-fold reduction in the sampling rate. The audio demo is available at https://qiangchunyu.github.io/VQCTAP/.
@@ I Introduction
Contrastive learning has proven to be an effective method for cross-modal modeling. It works by differentiating target samples (positive) from distractor samples (negatives) based on an anchor representation. The objective is to maximize the similarity between the anchor and positive samples while minimizing the similarity between the anchor and negative samples. This approach has been widely applied in text-image cross-modal representation, with notable examples including OpenAI’s CLIP[1], Florence[2], and ALIGN[3]. In the field of text-audio cross-modal representation, CLIP-based models have also been developed, including Wav2CLIP[4], AudioCLIP[5], and CLAP[6]. In the field of multi-modal representation, pre-trained representation models have emerged that use a single modality as a link to connect other modalities. For example, ImageBind[7] achieves joint modeling of multiple modalities based on the image modality, while LanguageBind[8] achieves joint modeling of multiple modalities based on the language modality. These models focus on extracting global descriptive information; for instance, CLAP aims to enhance performance in downstream audio classification tasks. However, their emphasis on global information limits their applicability to tasks requiring fine-grained sequence generation and recognition.
@@
As shown in Figure 1, the speech processing tasks of TTS and ASR are typical cross-modal applications. For tasks such as TTS, VC and ASR, a fine-grained (frame-level) sequence cross-modal representation(semantic coding) between text-speech modalities is expected. The intermediate representation extracted from speech should serve as a ”bridge” between text and acoustic information. It should emphasize linguistic content while de-emphasizing paralinguistic information such as speaker identity and acoustic details. Self-supervised representation learning methods, such as Wav2Vec2.0[9], Wav2Vec-C[10], VQ-Wav2Vec[11], HuBERT[12], and W2V-BERT[13], have become mainstream in the speech representation field. These methods offer the prospect of a universal model that can benefit a wide range of tasks and domains. However, when applied to VC and TTS tasks, such as SPEAR-TTS[14], these methods encounter issues of redundancy and dimensionality explosion. This is because the lack of explicit text supervision in these methods results in learned representations that are not fully decoupled semantic information. Supervised frame-level representation learning methods aim to explicitly introduce text information as supervision. For example, Phonetic Posteriorgrams (PPGs)[15] are calculated based on ASR acoustic models. Although PPGs have been widely used in various speech processing tasks, they are essentially text information, and the frame-level features are independent of each other, lacking the integration of contextual semantic information. Moreover, these methods rely on expensive annotated text-speech pairs. In this case, the model’s representational capacity is limited to the distribution of the text-speech paired data, resulting in poor robustness for unlabeled speech-only data (e.g., different styles, environmental sounds, channels, or qualities).
@@
In our previous three works: 1) We identified issues of information redundancy and dimensional explosion in existing semantic coding methods[16]. 2) We proposed a semantic coding method called ”Contrastive Token-Acoustic Pretraining (CTAP)[17].” 3) Building on CTAP, we developed a minimally-supervised high-fidelity speech synthesis method[18]. Building on these foundational works, we now introduce VQ-CTAP, which uses the cross-modal aligned sequence transcoder to bring text and speech into a joint multimodal space, facilitating the connection between text and speech at the frame level. Our contributions in this paper are as follows:
@@ II-B Contrastive Learning
In the audio field, CLIP-based models have also been developed, including Wav2CLIP[4], AudioCLIP[5], and CLAP[6]. These models focus on extracting global descriptive information from audio, with the primary goal of improving the performance of downstream audio classification tasks. However, their emphasis on global information limits their applicability to fine-grained generation and recognition tasks such as TTS, VC, and ASR, which require frame-level representations.
@@ II-C TTS, VC, and ASR tasks
The ASR is the inverse task of TTS, aims to predict text from speech. ASR is moving towards speech representation learning to make full use of the large-scale unsupervised speech datasets[26, 12, 23, 24]. The MLM[21, 9] based self-supervised learning machanism learn discrete masked speech prediction to derive the augmented speech representation, ASR plays as the downstream speech task benefits from the strong speech representation from large-scale datas based self-supervised learning method[21, 22, 23]
@@ III Method
As illustrated in Figure 2, the main body of VQ-CTAP is the cross-modal aligned sequence transcoder, which comprises five main components: a speech encoder, a phoneme encoder, a prompt encoder, a phoneme decoder, and a speech decoder. A length regulator is employed to address the mismatch between the lengths of phoneme sequences and speech sequences. The input consists of text-speech pairs and unlabeled random speech from outside the target domain. The phoneme and speech representations are concatenated in a joint multimodal space. A contrastive learning approach is used to learn this space by capturing the frame-level (dis)similarity between speech and phoneme pairs. The speech encoder produces discrete embedding at a rate of 25Hz from the 24kHz input waveform, which is a 960-fold reduction in the sampling rate. Vector quantization further removes paralinguistic information irrelevant to semantics from the speech representations. The two decoders adapt the learned representations for downstream tasks such as TTS, VC, and ASR. The method supports both supervised and unsupervised training. To enable domain-enhanced representation learning, semantic-transfer-wise paralinguistic consistency loss is introduced, computed using the unlabeled random speech from outside the target domain. The pre-trained VQ-CTAP can be directly applied to TTS, VC, and ASR tasks without fine-tuning, exhibiting a plug-and-play capability.
@@ III-A Cross-Modal Aligned Sequence Transcoder
Encoders: As illustrated in Figure 2 (a), $S_{in}$ denotes the input batch of target speech data, $S_{in}\in\mathbb{R}^{B\times T_{s}\times D_{s}}$, where $B$ is the batch size, $T_{s}$ is the number of time frames, and $D_{s}$ is the number of spectral components (mel-spectrogram bands). $P_{in}$ denotes the input batch of phoneme data, $P_{in}\in\mathbb{R}^{B\times T_{p}\times D_{p}}$, where $T_{p}$ is the number of phoneme codes and $D_{p}$ is the dimensionality of the phoneme codes. During training, the ground-truth duration is used to extend the length of the phoneme sequences to $T_{s}$. During inference, the corresponding predicted duration is used. After the length regulator, the phoneme sequences become $P_{in}\in\mathbb{R}^{B\times T_{s}\times D_{p}}$, having the same length as the target speech sequences $S_{in}$. The unlabeled random speech $R_{in}$ from outside the target domain is randomly selected, where $R_{in}\in\mathbb{R}^{B\times T_{r}\times D_{s}}$. The batch size $B$ and the number of spectral components $D_{s}$ are consistent with those of $S_{in}$, while typically $T_{r}\neq T_{s}$. The $P_{in}$, $S_{in}$ and $R_{in}$ are then passed through a phoneme encoder and a speech encoder, respectively:
$P$ $={PhonemeEncoder}(P_{in});$ (1)
$S$ $={SpeechEncoder}(S_{in});$
$R$ $={SpeechEncoder}(R_{in})$
@@
${G}=PromptEncoder(G_{in})$ (2)
$G$ denotes the prompt paralinguistic embedding, where $G$ is a fixed-length global vector. $G\in\mathbb{R}^{B\times D_{g}}$, where $D_{g}$ is the number of dimensions. The prompt encoder is a VAE-based model [46] that extracts paralinguistic information to address the one-to-many problem in TTS and VC tasks, such as timbre, style, and prosody, from the prompt speech. To address the KL collapse problem[47], a margin $\Delta$ is introduced to limit the minimum value of the KL loss as shown:
$\mathcal{L}_{kl}=max(0,D_{KL}[\mathcal{N}({\mu},{\sigma}^{2})||\mathcal{N}(0,I)]-\Delta)$ (3)
Decoders: As shown in Figure 2 (a), a discrete codebook component is added to the network for vector quantization processing, which further removes semantic-irrelevant paralinguistic information from the speech representation. $S$ and $R$ are compared with all vectors in the codebook, and the codebook vectors with the closest Euclidean distance are used as the vector-quantized speech representation $\hat{S}$ and $\hat{R}$. The vector quantization loss consists of two parts: the commitment loss (which makes the speech representation conform to its closest codebook vector as much as possible) and the codebook loss (which makes the selected codebook vector as close as possible to the speech representation). $sg[\cdot]$ represents ”stop gradient”, and $\boldsymbol{e}$ is the codebook vector.
@@
To utilize the pretrained speech representation for downstream ASR task, $S$ is input into a phoneme decoder to predict the phoneme sequence:
$P_{s}$ $=PhonemeDecoder(\hat{S})$ (5)
@@
$S_{s}$ $=SpeechDecoder(\hat{S},G);$ (7)
$R_{r}$ $=SpeechDecoder(\hat{R},G)$
@@ III-B Semantic-Transfer-Wise Paralinguistic Consistency Loss
Most existing methods use paired tuples of ground truth and synthesized features to calculate the cycle consistency loss. However, due to factors such as ”teacher forcing” and MSE loss, these two features are nearly identical during the training phase, resulting in suboptimal performance. During training, the combination of paralinguistic embedding and transferred semantic embedding (e.g., $G$ from speaker 1 and $\hat{R}$ from speaker 2) as input poses an unseen problem, since there is no ground truth to compute the reconstruction loss. Existing cross-modal fine-grained (frame-level) pre-training methods that aim to explicitly introduce text information as supervision rely on expensive annotated text-speech pairs. In this case, the model’s representational capacity is limited to the distribution of the text-speech paired data, resulting in poor robustness for unlabeled data (e.g., different styles, environmental sounds, channels, or qualities).
@@
III-C Token-Acoustic Contrastive Loss
With $S_{re}$ and $P_{re}$ now comparable, we can measure their similarity:
$C=\tau*({S}_{re}\cdot{P}_{re}^{\top})$ (10)
The contrastive loss is calculated:
where
@@ IV-A TTS Pipeline
Figure 4: The architecture of sequence-aware semantic connector
@@ IV-A1 Sequence-Aware Semantic Connector
Figure 4 illustrates the architecture of the sequence-aware semantic connector, which employs a bidirectional dilated convolution structure with $N$ residual layers organized into $m$ blocks, each containing $n=\frac{N}{m}$ layers. Within each block, the dilation is doubled at each layer. Skip connections from all residual layers are summed, akin to the approach used in WaveNet [51]. The model incorporates phoneme embedding as conditional information, which is processed by a transformer encoder and added as bias terms to the dilated convolution in each residual layer. The diffusion step embedding is broadcast over length and added to the input of each residual layer.
@@ IV-B VC Pipeline
VQ-CTAP can be directly applied to the VC task without fine-tuning, as illustrated in Figure 3(b). In this process, the weights of the prompt encoder, speech encoder, and speech decoder remain frozen. The pre-trained speech encoder extracts the speech embedding of the source speaker, while the pre-trained prompt encoder extracts the prompt paralinguistic embedding of the target speaker. The speech embedding $S$ is then quantized using a codebook to obtain the quantized speech embedding $\hat{S}$. Subsequently, $\hat{S}$ and prompt paralinguistic embedding is input together into the pre-trained speech decoder. The predicted mel spectrogram from the speech decoder combines the semantic content of the source speaker with the timbre of the target speaker.

### v2

@@ Abstract
Deep learning has brought significant improvements to the field of cross-modal representation learning. For tasks such as text-to-speech (TTS), voice conversion (VC), and automatic speech recognition (ASR), a cross-modal fine-grained (frame-level) sequence representation is desired, emphasizing the semantic content of the text modality while de-emphasizing the paralinguistic information of the speech modality. We propose a method called “Vector Quantized Contrastive Token-Acoustic Pre-training (VQ-CTAP)”, which uses the cross-modal aligned sequence transcoder to bring text and speech into a joint multimodal space, learning how to connect text and speech at the frame level. The proposed VQ-CTAP is a paradigm for cross-modal sequence representation learning, offering a promising solution for fine-grained generation and recognition tasks in speech processing. The VQ-CTAP can be directly applied to VC and ASR tasks without fine-tuning or additional structures. We propose a sequence-aware semantic connector, which connects multiple frozen pre-trained modules for the TTS task, exhibiting a plug-and-play capability. We design a stepping optimization strategy to ensure effective model convergence by gradually injecting and adjusting the influence of various loss components. Furthermore, we propose a semantic-transfer-wise paralinguistic consistency loss to enhance representational capabilities, allowing the model to better generalize to unseen data and capture the nuances of paralinguistic information. In addition, VQ-CTAP achieves high-compression speech coding at a rate of 25Hz from 24kHz input waveforms, which is a 960-fold reduction in the sampling rate. The experimental results demonstrate that while VQ-CTAP outperforms baseline methods in TTS and VC tasks, its performance on ASR tasks is suboptimal. The audio demo is available at https://qiangchunyu.github.io/VQCTAP/.
@@ I Introduction
Contrastive learning has proven to be an effective method for cross-modal modeling. This approach has been widely applied in text-image cross-modal representation, with notable examples including OpenAI’s CLIP[1], Florence[2], and ALIGN[3]. In the field of text-audio cross-modal representation, CLIP-based models have also been developed, including Wav2CLIP[4], AudioCLIP[5], and CLAP[6]. In the field of multi-modal representation, pre-trained representation models have emerged that use a single modality as a link to connect other modalities. For example, ImageBind[7] achieves joint modeling of multiple modalities based on the image modality, while LanguageBind[8] achieves joint modeling of multiple modalities based on the language modality. These models focus on extracting global descriptive information; for instance, CLAP aims to enhance performance in downstream audio classification tasks. However, their emphasis on global information limits their applicability to tasks requiring fine-grained sequence generation and recognition.
@@
As shown in Figure 1, the speech processing tasks of TTS and ASR are typical cross-modal applications. For tasks such as TTS, VC and ASR, a fine-grained (frame-level) sequence cross-modal representation(semantic coding) between text-speech modalities is expected. Semantic should provide a representation of speech where the linguistic content— from phonetics to semantics—is salient, while paralinguistic information such as speaker identity and acoustic details are removed[9]. The intermediate representation extracted from speech should serve as a “bridge” between text and acoustic information. It should emphasize linguistic content while de-emphasizing paralinguistic information such as speaker identity and acoustic details. Self-supervised representation learning methods, such as Wav2Vec2.0[10], Wav2Vec-C[11], VQ-Wav2Vec[12], HuBERT[13], and W2V-BERT[14], have become mainstream in the speech representation field. These methods offer the prospect of a universal model that can benefit a wide range of tasks and domains. However, when applied to VC and TTS tasks, such as SPEAR-TTS[9], these methods encounter issues of redundancy and dimensionality explosion. This is because the lack of explicit text supervision in these methods results in learned representations that are not fully decoupled semantic information. Supervised frame-level representation learning methods aim to explicitly introduce text information as supervision. For example, Phonetic Posteriorgrams (PPGs)[15] are calculated based on ASR acoustic models. Although PPGs have been widely used in various speech processing tasks, they are essentially text information, and the frame-level features are independent of each other, lacking the integration of contextual semantic information. Moreover, these methods rely on expensive annotated text-speech pairs. In this case, the model’s representational capacity is limited to the distribution of the text-speech paired data, resulting in poor robustness for unlabeled speech-only data (e.g., different styles, environmental sounds, channels, or qualities).
@@
In our previous three works: 1) We identified issues of information redundancy and dimensional explosion in existing semantic coding methods[16]. 2) We proposed a semantic coding method called “Contrastive Token-Acoustic Pretraining (CTAP)[17].” 3) Building on CTAP, we developed a minimally-supervised high-fidelity speech synthesis method[18]. Building on these foundational works, we now introduce VQ-CTAP, which uses the cross-modal aligned sequence transcoder to bring text and speech into a joint multimodal space, facilitating the connection between text and speech at the frame level. Our contributions in this paper are as follows:
@@ II-B Contrastive Learning
In the audio field, CLIP-based models have also been developed, including Wav2CLIP[4], AudioCLIP[5], and CLAP[6]. These models focus on extracting global descriptive information from audio, with the primary goal of improving the performance of downstream audio classification tasks. The downstream tasks applied in the experimental parts of these studies are primarily speech classification tasks, such as sound event classification, instrument classification, acoustic scene classification, emotion recognition, keyword spotting, vocal sound classification, and speaker counting. Since the global features extracted by these methods lose temporal information, they cannot be converted back into frame-level acoustic features. As a result, they are unsuitable for tasks like TTS, VC, and ASR[27].
@@ II-C TTS, VC, and ASR tasks
The ASR is the process of converting spoken language into text. This process encompasses several key components, including feature extraction, acoustic modeling, and language modeling. ASR is moving towards speech representation learning to make full use of the large-scale unsupervised speech datasets[26, 13, 23, 24]. The MLM[21, 10] based self-supervised learning machanism learn discrete masked speech prediction to derive the augmented speech representation, ASR plays as the downstream speech task benefits from the strong speech representation from large-scale datas based self-supervised learning method[21, 22, 23]
@@ III Method
As illustrated in Figure 2(a), the main body of VQ-CTAP is the cross-modal aligned sequence transcoder, which comprises five main components: a speech encoder, a phoneme encoder, a prompt encoder, a phoneme decoder, and a speech decoder. A length regulator is employed to address the mismatch between the lengths of phoneme sequences and speech sequences. The input consists of text-speech pairs and unlabeled random speech from outside the target domain. The phoneme and speech representations are concatenated in a joint multimodal space. As shown in Figure 2(c), a contrastive learning approach is used to learn this space by capturing the frame-level (dis)similarity between speech and phoneme pairs. The speech encoder produces discrete embedding at a rate of 25Hz from the 24kHz input waveform, which is a 960-fold reduction in the sampling rate. Vector quantization further removes paralinguistic information irrelevant to semantics from the speech representations. The two decoders adapt the learned representations for downstream tasks such as TTS, VC, and ASR. The method supports both supervised and unsupervised training. To enable domain-enhanced representation learning, semantic-transfer-wise paralinguistic consistency loss is introduced, computed using the unlabeled random speech from outside the target domain, as shown in Figure 2(b). The pre-trained VQ-CTAP can be directly applied to TTS, VC, and ASR tasks without fine-tuning, exhibiting a plug-and-play capability.
@@ III-A Cross-Modal Aligned Sequence Transcoder
Encoders: As illustrated in Figure 2 (a), $S_{in}$ denotes the input batch of target speech data, $S_{in}\in\mathbb{R}^{B\times T_{s}\times D_{s}}$, where $B$ is the batch size, $T_{s}$ is the number of time frames, and $D_{s}$ is the number of spectral components (mel-spectrogram bands). $\tilde{P}_{in}$ denotes the input batch of phoneme data, $\tilde{P}_{in}\in\mathbb{R}^{B\times T_{p}\times D_{p}}$, where $T_{p}$ is the number of phoneme codes and $D_{p}$ is the dimensionality of the phoneme codes. During training, the ground-truth duration is used to extend the length of the phoneme sequences to $T_{s}$. During inference, the corresponding predicted duration is used. After the length regulator, the phoneme sequences become $P_{in}\in\mathbb{R}^{B\times T_{s}\times D_{p}}$, having the same length as the target speech sequences $S_{in}$. The unlabeled random speech $R_{in}$ from outside the target domain is randomly selected, where $R_{in}\in\mathbb{R}^{B\times T_{r}\times D_{s}}$. The batch size $B$ and the number of spectral components $D_{s}$ are consistent with those of $S_{in}$, while typically $T_{r}\neq T_{s}$. The $P_{in}$, $S_{in}$ and $R_{in}$ are then passed through a phoneme encoder and a speech encoder, respectively:
$P$ $=\mathrm{PhonemeEncoder}(P_{in});$ (1)
$S$ $=\mathrm{SpeechEncoder}(S_{in});$
$R$ $=\mathrm{SpeechEncoder}(R_{in})$
@@
${G}=\mathrm{PromptEncoder}(G_{in})$ (2)
$G$ denotes the prompt paralinguistic embedding, where $G$ is a fixed-length global vector. $G\in\mathbb{R}^{B\times D_{g}}$, where $D_{g}$ is the number of dimensions. The prompt encoder is a VAE-based model [47] that extracts paralinguistic information to address the one-to-many problem in TTS and VC tasks, such as timbre, style, and prosody, from the prompt speech. In Equation (3), $D_{KL}$ refers to the KL loss, $\mathcal{N}(\cdot)$ represents a Gaussian distribution, and $({\hat{\mu}},{\hat{\sigma}})$ denotes the $(mean,variance)$ of the prompt paralinguistic latent space distribution. To address the KL collapse problem[48], a margin $\Delta$ is introduced to limit the minimum value of the KL loss as shown:
$\mathcal{L}_{kl}=max(0,D_{KL}[\mathcal{N}({\hat{\mu}},{\hat{\sigma}}^{2})||\mathcal{N}(0,I)]-\Delta)$ (3)
Decoders: As shown in Figure 2 (a), a discrete codebook component is added to the network for vector quantization processing, which further removes semantic-irrelevant paralinguistic information from the speech representation. In deep learning, a codebook refers to a collection of representative vectors or codes used to quantize data, typically in tasks like vector quantization or feature encoding. It helps reduce dimensionality, improve model efficiency, and enhance performance by mapping high-dimensional inputs to a finite set of meaningful representations[49]. $S$ and $R$ are compared with all vectors in the codebook, and the codebook vectors with the closest Euclidean distance are used as the vector-quantized speech representation $\hat{S}$ and $\hat{R}$. The vector quantization loss consists of two parts: the commitment loss (which makes the speech representation conform to its closest codebook vector as much as possible) and the codebook loss (which makes the selected codebook vector as close as possible to the speech representation). $sg[\cdot]$ represents “stop gradient”, and $\boldsymbol{e}$ is the codebook vector.
@@
To utilize the pretrained speech representation for downstream ASR task, $\hat{S}$ is fed into a phoneme decoder to predict the phoneme sequence:
$P_{s}$ $=\mathrm{PhonemeDecoder}(\hat{S})$ (5)
@@
$S_{s}$ $=\mathrm{SpeechDecoder}(\hat{S},G);$ (7)
$R_{r}$ $=\mathrm{SpeechDecoder}(\hat{R},G)$
@@ III-B Semantic-Transfer-Wise Paralinguistic Consistency Loss
Most existing methods use paired tuples of ground truth and synthesized features to calculate the cycle consistency loss[50, 51, 52, 47]. However, due to factors such as “teacher forcing” and MSE loss, these two features are nearly identical during the training phase, resulting in suboptimal performance. During training, the combination of paralinguistic embedding and transferred semantic embedding (e.g., $G$ from speaker 1 and $\hat{R}$ from speaker 2) as input poses an unseen problem, since there is no ground truth to compute the reconstruction loss. Existing cross-modal fine-grained (frame-level) pre-training methods that aim to explicitly introduce text information as supervision rely on expensive annotated text-speech pairs. In this case, the model’s representational capacity is limited to the distribution of the text-speech paired data, resulting in poor robustness for unlabeled data (e.g., different styles, environmental sounds, channels, or qualities).
@@
III-C Token-Acoustic Contrastive Loss
With $S_{re}$ and $P_{re}$ now comparable, we can measure their similarity:
$C=\tau*({S}_{re}\cdot{P}_{re}^{\top})$ (10)
The contrastive loss is calculated:
where
@@ III-D Stepping Optimization Strategy
Figure 4: The architecture of sequence-aware semantic connector
@@ IV-A1 Sequence-Aware Semantic Connector
Figure 4 illustrates the architecture of the sequence-aware semantic connector, which employs a bidirectional dilated convolution structure with $N$ residual layers organized into $m$ blocks, each containing $n=\frac{N}{m}$ layers. Within each block, the dilation is doubled at each layer. Skip connections from all residual layers are summed, akin to the approach used in WaveNet [55]. The dashed arrow indicates that the output of the current layer serves as the input to the next residual layer. The model incorporates phoneme embedding as conditional information, which is processed by a transformer encoder and added as bias terms to the dilated convolution in each residual layer. The diffusion step embedding is broadcast over length and added to the input of each residual layer.
@@ IV-B VC Pipeline
VQ-CTAP can be directly applied to the VC task without fine-tuning, as illustrated in Figure 3(b). In this process, the weights of the prompt encoder, speech encoder, and speech decoder remain frozen. The pre-trained speech encoder extracts the speech embedding of the source speaker, while the pre-trained prompt encoder extracts the prompt paralinguistic embedding of the target speaker. The speech embedding $S$ is then quantized using a codebook to obtain the quantized speech embedding $\hat{S}$. Subsequently, $\hat{S}$ and prompt paralinguistic embedding is fed together into the pre-trained speech decoder. The predicted mel spectrogram from the speech decoder combines the semantic content of the source speaker with the timbre of the target speaker.